Professional softball pitchers can deliver a pitch which can reach the blazing speed of 100 mph (miles per hour). A local team has drafted an up-and-coming right handed pitcher who can consistently throw at 40.67 m/s (91.00 mph). Assuming the ball has a mass 0.1424 kg and has this speed just before a batter would make contact, how much kinetic energy does the ball have?

A) Kball = ____ J

How high would the ball need to be dropped from to attain the same energy (neglect air resistance)? (The acceleration due to gravity is 9.81 m/s2.)

B) hball = ____ J
Answer #1

Solution)

Part a)

Kinetic Energy of ball = 0.5*m*v^2 = 0.5*0.1424*40.67*40.67 = 117.76 J (Ans)

======

Part b)

height h, of the ball

We know, v = sqrt(2*g*h)

40.67^2 = 2*9.81*h

h= 40.67*40.67/(2*9.81) = 84.39 m (Ans)

=======